I have tried several times to answer these two questions and I can't seem to get it right. Your help would be greatly appreciated!!
1) A triangle has a base of length 2x + 3 and a height of x + 3 and has an area of 27 square units. Find the length of the base and height. (A= 1/2BH)
2) A triangle has a base of length 3x + 1 and a height of x + 6 and has an area of 45 square units. Find the length of base and height. (A= 1/2BH)

Answer by Alan3354(69443)

1) A triangle has a base of length 2x + 3 and a height of x + 3 and has an area of 27 square units. Find the length of the base and height. (A= 1/2BH)
A = (1/2)*(2x+3)*(x+3) = 27
(2x+3)*(x+3) = 54
2x^2 + 9x + 9 = 54
2x^2 + 9x - 45 = 0
(2x + 15)*(x - 3) = 0
x = 3 (ignore the -15)
--------------------
2) A triangle has a base of length 3x + 1 and a height of x + 6 and has an area of 45 square units. Find the length of base and height. (A= 1/2BH)
2A = (3x+1)*(x+6) = 90
It's 9 by 10, so x = 3